OpenStudy (anonymous):

tanx- sec^2x/tanx=-cotx verify the identity

OpenStudy (anonymous):

Multiply through by tanx and you'll have something that should look familiar.

OpenStudy (snowsurf):

\[\tan^2(x) +1 = \sec^2(x)\] We will sue this identity. \[-\cot(x) = \tan(x) - \sec^2(x)/\tan(x) \] \[-\cot(x) = [\tan^2(x) -sex^2(x)]/\tan(x) \] Use the identity above \[-\cot(x)= [\tan^2(x) - (\tan^2(x) - 1]/\tan(x)\] \[-\cot(x) = - 1/\tan(x)\] 1/tan(x) is cot(x) so -cot(x) = -cot(x)

OpenStudy (snowsurf):

That 1 should be positive then become negative when distributing the negative sign.